OpenStudy (anonymous):

x^2-4x=21 solve by completing the square

OpenStudy (anonymous):

\[x^2-4x=21\] half of 4 is 2, and 2 squared is 4 so write \[(x-2)^2=21+4\] \[(x-2)^2=25\] therefore \(x-2=5\) or \(x-2=-5\) first equation gives \(x=7\) second gives \(x=-3\)

OpenStudy (anonymous):

my group thanks you!!!

OpenStudy (anonymous):

2012Sorching summer comes, in order to thank everyone, characteristic, novel style, varieties, low price and good quality, and the low sale price. Thank everyone ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== $33 True Religion jeans, Ed Hardy jeans,LV,Coogi jeans,Affliction jeans $30 Air Jordan shoes,Shox shoes,Gucci,LV shoes 50%Discount winter fashion :Sandle,t-shirt,caps,jerseys,handbag and brand watches!!! $15 Ed Hardy ,LV ,Gucci Bikini $15 Polo, Ed Hardy, Gucci, LV, Lacoste T-shirts $25 Coach,Gucci,LV,Prada,Juicy,Chanel handbag, $10 Gucci,Ed Hardy sunglasses $9 New Era caps. give you the unexpected harvest ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.fullmalls.com ) ===== ==== ( http://www.scnshop.com ) =====